a spherical conductor has inner radii 5 cm and outer radii 10 cm. a point charge of 1C is placed at a distance 3 cm from the center of the sphere. an external field of 20N/C is also applied. if the net force on the charge is 15 N away from the centre, find the force that the charge exerts on the sphere.
 

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
External applied field's effect is neutralized by charge induced on outer surface, i.e., their combined effect is zero. Hence, force applied by induced charges on outer surface should be equal and opposite to that applied by electric field. This is equal toqE=1×20=20NqE=1×20=20Ntoward left.15N force is only due to induced charged on the inside surface of the hollow sphere. Hence, net force by induced charges is15+20=35N